The banks in your area offer the following rates of interest on their savings accounts. If you want to open one of these accounts, which bank should you select?

Bank A: 1.845 percent APR with daily compounding.
Bank B: 1.840 percent APR with monthly compounding.
Bank C: 1.875 percent APR with annual compounding.
Bank D: 1.850 percent APR with quarterly compounding.
Bank E: 1.875 percent APR with semi-annual compounding.
A.

Bank D

B.

Bank E

C.

Bank B

D.

Bank C

E.

Bank A
Answer #1

EAR=(1+APR/m)^m-1
where m=compounding periods

Bank A:

EAR=(1+0.01845/365)^365-1

=1.862%(Approx).

Bank B:

EAR=(1+0.0184/12)^12-1

=1.856%(Approx).

Bank C:

EAR=(1+0.01875/1)^1-1

=1.875%

Bank D:

EAR=(1+0.0185/4)^4-1

=1.863%(Approx).

Bank E:

EAR=(1+0.01875/2)^2-1

=1.884%(Approx).

Hence Bank E must be selected having highest effective rate.